OpenStudy (frostarmy):

Will Fand and Medal Find the slope of the line that passes through the points (-1,4) and (2,-8). Question 2 options: A -12 B -6 C -4 D -4/3

OpenStudy (frostarmy):

Fan*

OpenStudy (coacoapuffprincess):

Ok

OpenStudy (coacoapuffprincess):

It wopuld be letter C

OpenStudy (coacoapuffprincess):

I think

563blackghost (563blackghost):

Use slope formula... \(\huge\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{ -8-4}{2+1}=slope }\)

OpenStudy (frostarmy):

@563blackghost i tried but overtime i use that i never seem to get the right answer

OpenStudy (frostarmy):

every time*

563blackghost (563blackghost):

You have to be careful on how you place your points. You might have flipped or forgot a sign. \(\huge\bf{-8-4=-12}\) \(\huge\bf{2+1=3}\) \(\huge\bf{\frac{-12}{3} \div \frac{3}{3} = -4}\)